# featuring a human-assisted proof

### v1

@@ 1 Introduction
Figure 1: An illustration of the perfect difference set $B=\{1,2,5,15,17\}\bmod{21}$. One can check that every difference from 11 to 2020 appears exactly once between the bold ticks; for example, $6$ is witnessed by $2-17\pmod{21}$. Suppose you have a favorite set, such as $A=\{1,5,15\}$. Erdős’s conjecture asks whether or not it is always possible to find a perfect difference set $B$ modulo some $v$ that contains your favorite set $A$, assuming of course that $A$ doesn’t already have a repeated difference. In this paper, we disprove this conjecture: you will be disappointed if your favorite set is $\{1,2,4,8,13\}$.
@@
The resulting formal proof is thousands of lines, nearly all of which were written by ChatGPT. Accordingly, we believe that ChatGPT is properly an author of the formal proof accompanying this paper [^2] However, the policy of the arXiv and other venues is that generative AI should not be listed as an author. Unfortunately, large language models are known to hallucinate and otherwise produce incorrect results, so we would not be able to trust this proof unless it was in a formal language. Therefore, we believe Lean is also an author, or perhaps ChatGPT and Lean may be credited together. Even so, while our formal proofs were written and verified by a large language model and a proof assistant, respectively, we stress that this paper was written the old-fashioned way by human authors.
@@ Acknowledgments
Our formalization would have been much more difficult without the proof assistant Lean. The authors are especially grateful to the Mathlib community for their unified library of mathematics in Lean4. Even with this extensive library, we found it surprisingly painful to successfully write proofs in Lean, so we also thank OpenAI for producing a large language model (namely, ChatGPT with GPT-5) that is capable of meeting a human mathematician halfway (or at least partway).
@@ 2 Some minor remarks on infinite sets
Does a Sidon set always extend
to a perfect difference set?
perfect difference set
@@
This result and its proof are stated formally by Hall [34, Theorem 3.1], but the proof is essentially a “just-do-it” [29, 30] greedy construction wrapped around the observation that there are only finitely many obstructions to adding a missing difference. ∎
@@ 5 Polarities and the remainder of Hall’s proof
$\{-8,-6,0,1,4\}$ doesn’t extend to a perfect difference set.
@@
$\{1,3,9,10,13\}$ is a Sidon set that doesn’t extend to a perfect difference set.
@@ 6 Lean statement
/-- A Sidon set ‘A‘ is a set where all pairwise sums ‘i + j‘ are unique,
up to swapping the addends. -/
@@
/-- ‘B‘ is a perfect difference set modulo ‘v‘ if there is a bijection between
non-zero residues mod ‘v‘ and distinct differences ‘a - b‘, where ‘a, b  B‘. -/
@@
  B.offDiag.BijOn (fun (a, b) => (a - b : ZMod v)) {x : ZMod v | x  0}
@@
Any finite Sidon set of natural numbers can be embedded in a perfect difference
set modulo ‘v‘ for some ‘v  0‘.
@@
The Sidon set {1, 2, 4, 8, 13} does not extend to a perfect difference set
modulo v for any nonnegative v.
@@
In the interest of reproducibility, we include here information about our versions of Lean [41] and the Mathlib [38] dependency:
Lean version4.24.0-rc1 919e297292280cdb27598edd4e03437be5850221
mathlib4               c69734131b8f3aff48af48197fe0480d6cc304cb
@@ 7 Discussion of the uses of artificial intelligence
Sadly, most of the rest of our actual experience with human-assisted proof wasn’t nearly this pleasant, though we are hopeful for a future that is closer to the above idealization. One of the early frustrations was while proving the next result, Proposition 19. It was promising at first, as ChatGPT itself constructed and proved the properties of the involution $p\mapsto p^{\prime}$ without any hints at all. However, it was completely stuck finishing the final step: that since the not-absolute points occur in pairs, there is an even number of them.
@@
Another, very simple but frequent issue that arose involved parity. Perhaps the ChatGPT model we used was trained heavily on mathlib3, the previous version of Mathlib, which has different definitions of basic concepts like even/odd and lemmas involving them. We were often unsuccessful with prompting ChatGPT to look up the appropriate definitions and use them. Thankfully this usually did not delay the work very much, as it’s a very simple matter. We mention it specifically because of the interaction between incompatible programming language version and large language models trained on the “wrong” version (from the perspective of the user).
Many of the issues we encountered would be alleviated greatly if the large language model interface we interacted with were integrated with Lean and reinforced appropriately with training in this interaction. We believe that even a small amount of such fine tuning would have made our specific “human-assisted proof” a much more pleasant interaction.

### v2

@@ 1 Introduction
Figure 1: An illustration of the perfect difference set $B=\{1,2,5,15,17\}\bmod{21}$. One can check that every difference from 11 to 2020 appears exactly once between the bold ticks; for example, $6$ is witnessed by $2-17\pmod{21}$. Suppose you have a favorite set, such as $A=\{1,5,15\}$. Erdős’s conjecture asks whether or not it is always possible to find a perfect difference set $B$ modulo some $v$ that contains your favorite set $A$, assuming of course that AA doesn’t already have a repeated difference. In this paper, we disprove this conjecture: you will be disappointed if your favorite set is $\{1,2,4,8,13\}$.
@@
The resulting formal proof is thousands of lines, nearly all of which were written by ChatGPT. Accordingly, we believe that ChatGPT is properly an author of the formal proof accompanying this paper [^2] However, the policy of the arXiv and other venues is that generative AI should not be listed as an author.. Unfortunately, large language models are known to hallucinate and otherwise produce incorrect results, so we would not be able to trust this proof unless it was in a formal language. Therefore, we believe Lean is also an author, or perhaps ChatGPT and Lean may be credited together. Even so, while our formal proofs were written and verified by a large language model and a proof assistant, respectively, we stress that this paper was written the old-fashioned way by human authors.
@@ Acknowledgments
Our formalization would have been much more difficult without the proof assistant Lean. The authors are especially grateful to the Mathlib community for their unified library of mathematics in Lean4, and to the Formal Conjectures authors for including this conjecture in their repository. Even with the extensive Mathlib library, we found it surprisingly painful to successfully write proofs in Lean, so we also thank OpenAI for producing a large language model (namely, ChatGPT with GPT-5 – but without ChatGPT Pro) that is capable of meeting a human mathematician halfway (or at least partway).
@@ 2 Some minor remarks on infinite sets
Does a Sidon set always extend to a perfect difference set? perfect difference set
@@
This result and its proof are stated formally by Hall [34, Theorem 3.1], but the proof is essentially a “just-do-it” [29, 30] greedy construction wrapped around the observation that there are only finitely many obstructions to adding a missing difference $d$ by inserting two elements $x$ and $x+d$. ∎
@@ 5 Polarities and the remainder of Hall’s proof
$\{-8,-6,0,1,4\}$ does not extend to a perfect difference set.
@@
$\{1,3,9,10,13\}$ is a Sidon set that does not extend to a perfect difference set.
@@ 6 Lean statement
/- This proof has been verified on Lean Toolchain
version leanprover/lean4:v4.24.0 and Mathlib version
f897ebcf72cd16f89ab4577d0c826cd14afaafc7 (v4.24.0) -/
@@
/-- A Sidon set ‘A‘ is a set where all pairwise sums
‘i + j‘ are unique, up to swapping the addends. -/
@@
/-- ‘B‘ is a perfect difference set modulo ‘v‘ if there
is a bijection between non-zero residues mod ‘v‘ and
 distinct differences ‘a - b‘, where ‘a, b  B‘. -/
@@
  B.offDiag.BijOn (fun (a, b) => (a - b : ZMod v))
  {x : ZMod v | x  0}
@@
Any finite Sidon set of natural numbers can be embedded
in a perfect difference set modulo ‘v‘ for some ‘v  0‘.
@@
The Sidon set {1, 2, 4, 8, 13} does not extend to a
perfect difference set modulo v for any nonnegative v.
@@
#print axioms not_erdos_707AM
-- ’not_erdos_707AM’ depends on axioms:
-- [propext, Classical.choice, Quot.sound]
@@ 7 Discussion of the uses of artificial intelligence
Sadly, most of the rest of our actual experience with human-assisted proof wasn’t nearly this pleasant (reflective of ChatGPT’s so-called “jagged intelligence”), though we are hopeful for a future that is closer to the above idealization. One of the early frustrations was while proving the next result, Proposition 19. It was promising at first, as ChatGPT itself constructed and proved the properties of the involution $p\mapsto p^{\prime}$ without any hints at all. However, it was completely stuck finishing the final step: that since the non-absolute points occur in pairs, there is an even number of them.
@@
Another, very simple but frequent issue that arose involved parity. Perhaps the ChatGPT model we used was trained heavily on mathlib3, the previous version of Mathlib, which has different definitions of basic concepts like even/odd and lemmas involving them. We were often unsuccessful with prompting ChatGPT to look up the appropriate definitions and use them. Thankfully this usually did not delay the work very much, as it’s a very simple matter. In general, it can be frustrating when different versions of a programming language are incompatible and the large language model of choice was trained on the “wrong” version, relative to the use case.
Many of the issues we encountered would be alleviated greatly if the large language model interface we interacted with were integrated with Lean and appropriately trained for this interaction. We believe that even a small amount of such fine tuning would have made our specific “human-assisted proof” a much more pleasant interaction.